I do not understand the steps in a few fractional equation problems, but I would like to know how to work this one out, so I wont be confused in the future. n - 16/n = 96/5
Answer by edjones(8007)

n - 16/n = 96/5
5n^2-80=96n multiply each side by 5n (LCM)
5n^2-96n-80=0
5n^2+4n-100n-80=0
n(5n+4)-20(5n+4)=0 factoring by grouping.
(n-20)(5n+4)=0
5n=-4, n=-4/5
n=20
.
Ed